OpenStudy (anonymous):

Solve f(x) = −4x − 1, find f(−2).

OpenStudy (anonymous):

f(-2) =-4(-2)-1

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

-4(-2)-1=8-1=7

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

-4(-2) =8 8-1=7

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

Uh not really...

OpenStudy (anonymous):

If you multiply two negative numbers, yo get a positive number.

OpenStudy (anonymous):

\[f(x)=-4x-1\]\[f(-2)=-4(-2)-1\]\[f(-2)=8-1\]\[f(-2)=7\] \[\large :)\]

OpenStudy (anonymous):

ohhhhhh ok i see i dident see the other negative lol ty im posting anothe can u guys help on that too?